Circular No. 8181

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002li–2002ls P. Ruiz-Lapuente, University of Barcelona, on behalf of the European

Supernova Cosmology Consortium (cf. IAUC 7406, plus A. Balastegui, G. Sainton, R. Pascoal, and R. Amanullah), reports the discovery of eleven supernovae found with the Isaac Newton 2.5-m telescope (INT) via the project “Ω and Λ from Supernovae, and the Physics of Supernova Explosions” (undertaken within the International Time Programme on the telescopes of the European Northern Observatory at La Palma):

SN  2002 UT  α_{2000}  δ_{2000}  g′  Offset
2002li June 4 15^{h}59^{m}03.^{s}08 +54^{o}18′16′′.0 20.6 0′′.1 W, 0′′.2 S
2002lj June 5 16 19 19.65 +53 09 54.2 19.7 0′′.2 W
2002lk June 6 16 06 55.92 +55 28 18.2 18.3 0′′.7 W, 0′′.2 N
2002ll June 6 16 37 08.30 +41 24 37.7 20.5 0′′.1 E, 0′′.1 N
2002lm June 6 16 39 04.19 +42 21 32.1 22.9 0′′.2 S
2002ln June 6 16 39 24.93 +41 47 29.0 22.5 1′′.1 W, 8′′.3 S
2002lo June 6 16 39 56.42 +42 19 20.5 21.9 0′′.6 E, 1′′.3 S
2002lp June 6 16 40 11.45 +42 28 30.2 20.4 0′′.2 E, 2′′.0 N
2002lq June 6 16 40 28.83 +41 14 09.1 21.9 4′′.5 E, 0′′.7 S
2002lr June 6 22 33 12.59 + 1 05 56.7 20.8 3′′.0 S
2002ls June 7 22 37 40.89 − 0 38 00.8 22.6 0′′.7 E, 1′′.0 S

All were discovered near maximum and all were confirmed with spectra obtained on June 10 and 11 at the William Herschel Telescope; followup confirmation and photometry were performed at the Jacobus Kapteyn Telescope, INT, Nordic Optical Telescope, and the Telescopio Nazionale Galileo. All objects were found to be type-Ia supernovae, except 2002lm (type-II), 2002ln (possible type-II), and 2002ll and 2002lr (possible type-Ia) — with the following measured redshifts: SN 2002li, 0.327; 2002lj, 0.14; 2002lk, 0.12; 2002ll, 0.05; 2002lm, 0.28; 2002ln, 0.14; 2002lo, 0.14; 2002lp, 0.114; 2002lq, 0.07?; 2002lr, 0.21; 2002ls, 0.25? Additional g′ magnitudes: 2002li, 1999 July 19 UT, [24.5; 2002 June 7.95, 20.5. 2002lj, 1999 Aug. 13, [24.9; 2002 June 7.89, 19.8. 2002lk, 1999 Aug. 8, [24.8; 2002 June 7.92, 18.1. 2002ll, 1999 Aug. 15, [24.7; 2002 June 8.11, 20.6. 2002lm, 2000 June 26, [24.7; 2002 June 8.00, 23.0. 2002ln, 2000 June 26, [24.3. 2002lo, 2000 June 27, [24.6; 2002 June 7.00, 21.8. 2002lp, 2000 June 27, [24.7; 2002 June 7.00, 20.4. 2002lq, 1999 Aug. 17, [24.8. 2002lr, 1999 July 22, [24.6; 2002 June 8.17, 21.0. 2002ls, 1999 July 17, [24.5; 2002 June 8.20, 22.6.

2003 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green